OpenStudy (anonymous):

The angle of elevation to the top of a building is found to be 13 ∘ from the ground at a distance of 5000 feet from the base of the building. Find the height of the building.

OpenStudy (anonymous):

 tan(13)=x/5000 x=5000(tan(13))=______ feet